OpenStudy (anonymous):

what is the value of the discriminant in the equation shown below? x^2-6x+4=0

OpenStudy (anonymous):

36-16=20

OpenStudy (anonymous):

how did you get that?

OpenStudy (amistre64):

what is a discriminate?

OpenStudy (anonymous):

The term underneath the radical in the quadratic equation

OpenStudy (amistre64):

good; and what is the general set up for that term?

OpenStudy (amistre64):

by general i mean the "letter" version

OpenStudy (anonymous):

square root of b^2-4ac...?

OpenStudy (amistre64):

yep; but we can omit the square root symbol ...

OpenStudy (amistre64):

b^2 - 4ac would be good

OpenStudy (amistre64):

what is the general set up for any given quadratic equation?

OpenStudy (anonymous):

x=-b+or - square root of b^2-4ac/2a

OpenStudy (amistre64):

.... thats good, not quite what i had in mind, but good :) the general set up of a quadratic equation is: ax^2 +bx + c the specific set up for this equation is: x^2 -6x +4 from this; what do we determine a b and c to be?

OpenStudy (amistre64):

maybe this might be easier to relate them [a] x^2 [+b] x [+c] [1] x^2 [-6] x [+4]

OpenStudy (anonymous):

so that is how hashir got 36-16=20

OpenStudy (amistre64):

yes :) (-6)^2 -4(1)(4) 36 - 4(4) 36 - 16